OpenStudy (anonymous):

WILL MEDAL AND FAN geometry(quick) Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 -2x + y2 - 6y = 26

OpenStudy (welshfella):

first complete the squares (x -1)^2 - 1 + (y -3)^2 - 9 = 26 (x -1)^2 + (y - 3)^2 = 36 this is the general form for a circle you can pick out the center and radius from this

OpenStudy (anonymous):

Thank you!

OpenStudy (welshfella):

yw